Circular No. 7694

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q3 (SOHO) Further to IAUC 7689, D. Hammer reports astrometric measurements

of a comet found by Chen Dong Hua on SOHO website images. The object was first detected (at the position given below) in LASCO C2 coronagraph images, moving southward from directly underneath the occulter; the comet brightened and showed a nice tail before entering the C3 field-of-view, where it began to fade, as indicated by the following additional V magnitudes provided by D. Biesecker and Hammer: Aug. 25.393 UT, 7.6; 25.463, 7.4; 25.977, 5.7; 26.102, 4.8; 26.221, 5.2; 26.227, 4.6; 26.446, 5.8; 26.811, 8.0.

2001 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 25.338  10^{h}15.^{m}1  +12^{o}35′  8.3

The reduced positions, the following orbital elements, and an ephemeris by B. G. Marsden appear on MPEC 2001-Q60.

T = 2001 Aug. 26.51 TT  ω = 268.^{o}11
Ω = 155.03 }2000.0
q = 0.0373 AU  i = 106.63

V4643 SAGITTARII N. M. Ashok and D. P. K. Banerjee, Physical Research Laboratory,

Ahmedabad; and W. P. Varricatt, Joint Astronomy Centre, Hilo, report spectroscopy of V4643 Sgr at 1.05–2.55 μm with the U.K. Infrared Telescope (+ CGS4) in June and on Aug. 12: “The K-band spectrum taken on June 16 is dominated by the Si VI and VII lines at 1.963 and 2.476 μm, respectively, indicating that the nova has entered the coronal phase. These lines show multiple components and large expansion velocity (FWZI = 9000–10000 km/s). The Ca VIII line at 2.323 μm is also detected. The J-band spectrum taken on June 28 and the H-band spectrum taken on June 29 show that the H I emission lines at Paβ, Paγ, and Br_{10} to Br_{14} have considerably weakened, and the O I line at 1.128 μm has disappeared (having been present in the Mt. Abu spectra taken on Mar. 3; IAUC 7599). The J-, H-, and K-band spectra of Aug. 12 are similar to those of June, except that the Si VI and VII lines have weakened considerably, and Ca VIII line has faded below detection limit.”

2001 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green